The High Five Beanbag company can manufactrue a certain number of beanbags per day for $180. If they manufacture thirty more beanbags per day, the manufacturing price drops be $.50 per unit while the total daily manufacturing cost of $180 remains the same. How many beanbags did the company originally manufacture per day?
Answer by stanbon(48568)

The High Five Beanbag company can manufactrue a certain number of beanbags per day for $180. If they manufacture thirty more beanbags per day, the manufacturing price drops be $.50 per unit while the total daily manufacturing cost of $180 remains the same. How many beanbags did the company originally manufacture per day?
Original Price Data:
# of bags = x; total cost= $180; price per bag= 180/x dollars
New Price Data:
# of bags = x+30; cost=$180; price per bag = 180/(x+30) dollars
EQUATION:
Original price per bag = New price per bag + 1/2 dollar
180/x = 180/(x+30) + 1/2
Multiply thru by 2x(x+30) to get:
2(180)(x+30) =2(180)(x) +x(x+30)
360x+(60)(180)=360x+x^2+30x
x^2+30x-10800=0
(x-90)(x+120)=0
x=90 or x=-120
Only x=90 is applicable
Original number of bags produced is 90.
Cheers,
Stan H.